Lydia decides to mix grades of gasoline in her truck. She puts in 6 gallons of regular and 9 gallons of premium for a total cost of $56.64. If premium gasoline costs $0.21 more per gallon than regular, what was the price of each grade of gasoline?
Answer by dragonwalker(73)

Let us call the cost of regular 'g' so the cost of premium is 'g+0.21'
Multiply the number of gallons by its cost per gallon for each type and add together which will equal $56.64 so:
6g + 9(g+0.21)=56.64
solve for g:
6g + 9g + 1.89 = 56.64
15g = 56.64-1.89 (when moving a number from one side to the other side of an equation change the sign)
15g = 54.75
divide both sides by 15 to find g:
15g/15 = 54.75/15
g = 3.65
So the cost of regular is $3.65 and therefore the cost of premium is g+0.21 = $3.86
To check:
6(3.65) + 9(3.86) = 56.64 !!!